# Genie: Smart ROS-based Caching for Connected Autonomous Robots

### v1

@@ V Related Work
Caching and Intelligence Caching. The idea of reusing computation to reduce cost and computation overhead is not new. FoggyCache proposed a technique to help close-quarters devices re-use redundant computations [18, 19]. Nonetheless, such techniques rely on a unified server [20] and do not provide driving-specific benefits. Similar approaches have been proposed for the cloud [21, 22], but rely on a central caching API [23, 24] to build and distribute the cache. Such techniques can be also applied to timing-sensitive AR/VR applications [25], storage systems [26], mobile applications [27, 28], robotic applications [29].
Collaborative Sensing. Several approaches employ collaborative sensing [30, 31], such as Potluck [32], but focus on single devices. Others such as Darwin phones [33] rely on a central unit to distribute an intelligence task, which is different from the problem addressed in this paper.
Real-Time Autonomous Driving. Reusability, especially in the context of distributed systems is not generally used in Real-Time Systems. However, caching and similar techniques can be an invaluable tool to meet latency requirements. Others have sought to improve latency on embedded platforms via various techniques such as approximation [34], smart scheduling [35, 36, 37, 38], and memory management [39, 40]. Nonetheless, the concept of caching for autonomous vehicles is unique because it exposes new challenges such as non-intrusive cache management and collaborative cache.
@@ VI Conclusion
In this paper, we identify and address three key challenges, raised by an industry partner, that hinder practical implementation of edge-computing for connected autonomous vehicles. We develop Genie as an answer to these challenges which is proven to be effective via implementation and evaluation using realistic workloads and the latest relevant platform. Efforts to expand Genie to applications beyond autonomous driving, as well as a port to ROS 2, are underway.

### v2

@@ V Related Work
Caching and Intelligent Caching. The concept of reusing computation to minimize costs is well-established. Techniques like FoggyCache enable close-quarters devices to reuse redundant computations [19, 20], but depend on centralized servers and lack driving-specific benefits. Similar methods exist for cloud environments [21, 22], relying on centralized APIs [23, 24] to manage caching, applicable to AR/VR [25], storage systems [26], mobile apps [27, 28], and robotics [29]. A few concurrent works like FogROS [30], FogROS2 [31], and Schafhalter et. al [32] also focus improving latency performance in robots by using cloud hardware.
AI-based Collaborative Sensing. AI is expected to make great impacts on various robotic fields, including home services [33], healthcare [34, 35, 36, 37], and transportation [38, 39, 40, 41], etc., in 2030 according to Stanford’s report [42]. Deep learning approaches have become prominent in autonomous driving [43, 44]. Approaches like collaborative sensing [45, 46] focus on individual devices or central units for task distribution, as seen in Potluck [47] and Darwin phones [48]. This contrasts with our decentralized approach. EMP [49] and AutoCast [50] proposed collaborative perception leveraging the edge to improve overall accuracy via decentralized data sharing among vehicles. Genie has the potential to integrate these solutions in ROS-based scenarios.
Real-Time Autonomous Driving. Reusability in distributed systems is uncommon in real-time contexts. Caching offers significant potential for meeting latency requirements, as seen in embedded platforms using approximation [51], scheduling [52, 53, 54, 55], memory management [56, 57], and software-hardware co-design [58]. Our solution Genie focuses on autonomous vehicle caching and introduces unique challenges, such as non-intrusive cache management and collaborative caching. Furthermore, we plan to adapt Genie to more safety-critical multi-robot scenarios [59, 60, 61, 62, 63].
@@ VI Conclusion
This paper addresses three key challenges identified by an industry partner that impede the practical implementation of edge computing for connected autonomous vehicles. We developed Genie, demonstrating its effectiveness through implementation and evaluation with realistic workloads and contemporary platforms. Future work aims to extend Genie beyond autonomous driving and adapt it to ROS 2 for more modern robotic system solutions.